OpenStudy (anonymous):

What is the solution to 8x-2=x^2?

OpenStudy (tkhunny):

Why not write it in standard Quadratic Equation Form and solve with your usual methods?

OpenStudy (anonymous):

I made the equation equal 0 and Im factoring.

OpenStudy (tkhunny):

It's not likely that it will factor. Quadratic Formula?

OpenStudy (anonymous):

The equation when equaling 0 is \[-x^2+8x-2=0\]

OpenStudy (tylerd):

.you could use completeing the square to factor i think.

OpenStudy (anonymous):

Like square both sides before making the equation equal 0?

OpenStudy (tkhunny):

Don't make your life difficult. Use \(x^{2} - 8x + 2 = 0\). This is standard Form. Completing the Square is the same as the quadratic formula. Unique answers don't care how you find the,.

OpenStudy (tylerd):

\[-x^2+8x-2=0\] \[-1(-x^2+8x-2)=0(-1)\] \[x^2-8x+2=0\] \[x^2-8x=-2\] \[x^2-8x+16=-2+16=14\] \[(x-4)(x-4)=14\] \[x-4= \pm \sqrt{14}\]

OpenStudy (tylerd):

quadratic formula is easier, I just dont think you should get to dependent on the quadratic formula.

OpenStudy (tkhunny):

Quadratic Formula and Completing the Square are the same thing in an algebra problem. In the calculus, you will be very glad that you are good at Completing the Square. You'll need that.

OpenStudy (tylerd):

Quadratic Formula \[ax^2+bx+c\] \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

I was with you until the 5th step. How did you get the 16=-2+16=14?

OpenStudy (tylerd):

ok so this is completeing the square method, and the 16 is we take half of 8 = 4 then square it to get 16. 8 is the "b" term

OpenStudy (tylerd):

er b coefficient

OpenStudy (tylerd):

half of b squared , then add that number to both sides.

OpenStudy (tylerd):

thats just how you get a perfect square so it factors out.

OpenStudy (tkhunny):

Think really hard on this concept: \((a+b)^{2} = a^{2} + 2ab + b^{2}\)

OpenStudy (tylerd):

if you just need to get the problem done fast, just plug it in to the quadratic formula.

OpenStudy (anonymous):

Ok, thank you.

OpenStudy (tylerd):

i think i got around x=7.74 and x=0.25 something around there.